Session: Nov 1999 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Cattani at Sat Nov 13 15:57:35 MET 1999) Success

Torun (K.Borkowski at Sat Nov 13 16:57:37 MET 1999) Success
No known problems.

Shanghai (Huang Xinyong at Mon Nov 15 01:07:41 MET 1999) Success

Effelsberg (A. Kraus at Mon Nov 15 09:13:42 MET 1999) Success
By mistake a thin tape instead of a thick one was used.

Noto (C. Trigilio at Mon Nov 15 09:57:56 MET 1999) Success
Cloudy sky, no known problem.

Onsala (A. Polatidis at Mon Nov 15 10:37:49 MET 1999) Success
No known problems. Clear skies

Westerbork (Tony Foley at Mon Nov 15 15:04:24 MET 1999) Success
12 telescopes in array No known problems

Jodrell (Alastair Gunn at Tue Nov 16 11:13:10 MET 1999) Success
Extreme RFI. No known problems.

Cambridge (Alastair Gunn at Tue Nov 16 12:13:44 MET 1999) Minor failures
Problems reading procedure file at start of experiment. Data good from about 08:10 UT.

Westerbork (Tony Foley at Wed Nov 17 13:22:22 MET 1999) Minor failures
Problem discovered late; a previously used tape (USN00338) which had tracks recorded at -319 and +31 micron was used without erasing.This will probably cause passes 1,2,15 and 16 to be bad.

Urumqi (Liu Xiang at Tue Dec 7 06:08:22 MET 1999) Success
No known problem.

Calibration (Jean\-Francois Desmurs at Wed Dec 15 12:47:02 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov99 directory under the name: es027.antab. Read the comments in the file to get further information.